OpenStudy (anonymous):

write the equation for the transformation of the graph y=f(x). *translated 5 units right, reflected across the x-axis*

OpenStudy (anonymous):

a horizontal translation will move the graph left or right on the and the change happens directly to the x so if you move it to the right 6 units you would right f(x-6) or to the left 3 units f(x+3) notice how if you want it to go to the left you use a "+" sign and right s "-" sign. for a reflection across the x-axis you add a "-" sign directly in front of the f... y=-f(x). for a reflection across the y-axis the "-" sign go directly in front of the x... y=f(-x) y=-f(x-5)

OpenStudy (anonymous):

Can you help me with one more? @Jared25

OpenStudy (anonymous):

Yeah Sure

OpenStudy (anonymous):

g(x)=f(x)-4 and also Vertex (-3,0)

OpenStudy (anonymous):

what exactly does the problem ask you to do?

OpenStudy (anonymous):

The first one is: Describe the transformation(s) of the parent function f(x) The second one is: Write the equation for each translation of the graph y=|x|

OpenStudy (anonymous):

Okay for the first: For a horizontal translation Then change in the function happens like f(x) to f(x-5) you would say that the graph shifted 5 units to the right. the change happens in the parentheses. For a Vertical translation the change happens outside of the parentheses so f(x) to f(x)+5 would be a vertical translation and you would say that the graph has been shifted up 5 units. So for your first problem you would say something like "the graph has shifted down 4 unit"s. make sense?

OpenStudy (anonymous):

Yes @Jared25

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

did the vertex (-3,0) have to do with the first problem or a different problem?

OpenStudy (anonymous):

I think Its a different one @Jared25

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

did you need help with anything else?

OpenStudy (anonymous):

@Bdavis15

OpenStudy (anonymous):

Yes @Jared25

OpenStudy (anonymous):

okay what with?

OpenStudy (anonymous):

Write the equation for each translation of the graph y=|x| Vertex 4,1

OpenStudy (anonymous):

not quite sure if I'm understand this one correctly. but if i am then it would be y=|x-4|+1

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

no prob